OpenStudy (anonymous):

Find the exact value of tan 75 degrees

OpenStudy (anonymous):

use the fact that \(75=30+45\) and then the "addition angle" formula for tangent

OpenStudy (anonymous):

does that help at all, or would you like me to walk you through it? use \[\tan(\alpha+\beta)=\frac{\tan(\alpha)+\tan(\beta)}{1-\tan(\alpha)\tan(\beta)}\] with \[\alpha =45, \beta = 30\]

OpenStudy (anonymous):

Yes thanks